OpenStudy (anonymous):

What is the slope of a line parallel to 4x + y = -1?

OpenStudy (anonymous):

A line parallel to 4x + y = -1 will have the same slope as the line 4x + y = -1.

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

Do you know how to find the slope of the line 4x + y = -1?

OpenStudy (anonymous):

no I do not.

OpenStudy (anonymous):

The genral equation of a line is y = mx + b, where m is the slope. So we will take our equation and solve for y: 4x + y = -1 Subtract 4x from both sides, we get y = -4x - 1 so the slope is -4, the coefficient of the x-term.

OpenStudy (anonymous):

So any line parallel to this line will also have a slope of -4.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

can you show me how to do this one?

OpenStudy (anonymous):

What is the slope of a line perpendicular to x - 5y = -10